OpenStudy (anonymous):

What is the complete factorization of the polynomial below? x3 + 3x2 - x - 3 A. (x + 1)(x - 1)(x + 3) B. (x - 1)(x - 1)(x - 3) C. (x - 1)(x - 1)(x + 3) D. (x + 1)(x - 1)(x - 3)

OpenStudy (alexandervonhumboldt2):

Factor out common terms x^2(x+3)−(x+3) (x+3)(x^2−1) now use the diffrence of squares: a^2-b^2=(a-b)*(a+b) you can factor x^2-1=x^2-1^2=?

OpenStudy (alexandervonhumboldt2):

do you know how to factor x^2-1?